OpenStudy (anonymous):

a graph of 2x - y = 4

OpenStudy (anonymous):

I need the graph

OpenStudy (alexwee123):

that's nice

OpenStudy (anonymous):

Try solving for Y. 2x-y=4 2x-y-4=0 2x-4=y It should be easier to graph y=2x-4